# Double strangeness pentaquarks and other exotic hadrons in the $\Xi_{b}\to\Xi\ J/\Psi\ \phi$ decay

### v1

@@ III.1 Light sector
We first discuss the results of the light channels of the pseudoscalar meson baryon interaction, where we employ the dimensional regularisation scheme to compute the loop function. The values of the unknown subtraction constants are determined by imposing the loop function to coincide, for a regularisation scale of $\mu=1000~{}{\rm MeV}$, with the loop cut-off function with $\Lambda=800~{}{\rm MeV}$ at the corresponding threshold (Eq. (16)). We chose this value as it corresponds to the mass of the interchanged vector meson on the t-channel diagram which is integrated out when we take the $t\to 0$ limit. We will refer to this procedure as Model PB 1.
@@
Figure 3: The curves represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{3j}|$, obtained in Model PB1. The vertical coloured dashed lines represent the location of the thresholds of the different channels. The solid line shows the results when the heavy channels are neglected and the dashed one presents the results using all 9 the channels.
@@
Figure 4: Solid lines represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for a fixed channel $i$, obtained in Model PB 1. Subfigure (b) shows the region of the second pole with more details at higher scale. The vertical dashed lines represent the location of the thresholds of the different channels.
@@
$a_{i}$ $\Lambda$ $|g_{i}|$ $\chi_{i}$ $|g_{i}|$ $\chi_{i}$
$\pi\Xi(1456)$1.51 800 2.44 0.428 0.12 0.001
$\bar{K}\Lambda(1611)$1.32 800 1.91 0.169 0.27 0.006
$\bar{K}\Sigma(1689)$1.27 800 0.71 0.016 1.54 0.815
$\eta\Xi(1866)$1.47 800 0.49 0.004 1.06 0.029
@@
We can observe that the lowest generated resonance couples strongly to the $\pi\Xi$ and $\bar{K}\Lambda$ channels, which are open for decay, what actually explains its rather large width. In the PDG PDG one can find a state with similar mass, $\Xi(1620)$, which was seen decaying into the $\pi\Xi$ channel (Tab. 4). Even if our state is $50~{}{\rm MeV}$ lighter and its width is larger, we can see the potential of our model to generate a state in this region.
As it can also be seen in Table 3 our heavier state mostly couples to the $\bar{K}\Sigma$ and $\eta\Xi$ channels and it has a mass of $1685~{}{\rm MeV}$. According to PDG data there exists a $\Xi(1690)$ state at $1690\pm 10~{}{\rm MeV}$ with a width of $20\pm 15~{}{\rm MeV}$ (Tab. 4), which is compatible with our result. [^1] If this state were moved to a bit higher energy, the $\bar{K}\Sigma$ channel would be open for it to decay, and this would naturally lead to an increase of the width.
@@
Looking at the results, it is natural to think that we can modify the values of the subtraction constants within a reasonable range in order to accommodate the generated states to the experimental data. Therefore, we relax the condition for our loop function has to be equal to the cut-off loop function with $\Lambda=800~{}{\rm MeV}$ at the corresponding threshold Montana:2017kjw . By doing that we define a new model, called PB 2. The results of new calculations within Model PB2 are shown in Fig. 5 and in Table 3. We can see that our lighter resonance now appears at higher mass and it is much wider, while the heavier resonance is generated with a similar mass and with a slightly larger width. Looking at Table 4 we can see that now the mass of the light state is closer to the experimental energy of the $\Xi(1620)$. In this model, none of the equivalent cut-off values for the subtractions constants employed are smaller than $\Lambda=630~{}{\rm MeV}$, which is a reasonable lower limit value.
Figure 5: Solid lines represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for a fixed channel $i$, obtained in Model PB 2. The vertical dashed lines represent the location of the thresholds of the different channels.
Reviewing the literature Sekihara:2015qqa ; Kaiser:1995eg ; Oset:1997it ; Ramos:2002xh ; Garcia-Recio:2003ejq ; Gamermann:2011mq ; Nishibuchi:2023acl ; Nishibuchi:2022zfo ; Nishibuchi:2023nvi one can note that all the models based only on a leading interaction term in the potential, the so called Weinberg-Tomozawa term, see eq. (6), although can dynamically generate two poles, but the features of those could not reproduce the experimental masses and widths of $\Xi(1620)$ and $\Xi(1690)$ at the same time. Furthermore, the width of the $\Xi(1690)$ remains always rather small, around a few MeV. And our calculations do also show all these features.
In principle, the $\Xi(1620)$ resonance can be reproduced very well Nishibuchi:2023acl ; Nishibuchi:2022zfo ; Nishibuchi:2023nvi , but only allowing some of the subtracting constants to take unnatural-size values.
This limitations can be overcame, if one takes into consideration also the s- and u-channel Born diagrams and the NLO contribution, as it was done in Feijoo:2023wua ; Magas:2024mba ; Sarti:2023wlg , where both $\Xi(1620)$ and $\Xi(1690)$ are well reproduced by the model.
@@
Figure 6: Solid lines represent the sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB 1. The vertical coloured dashed lines represent the location of the thresholds of the different channels and the black one represents the mass position of the $\Xi(1950)$.
@@
$\rho\Xi(2089)$1.50 800 3.39 0.610 0.24 0.004
$\bar{K}^{*}\Lambda(2010)$1.42 800 1.82 0.358 0.52 0.0490
$\bar{K}^{*}\Sigma(2087)$1.46 800 0.98 0.047 2.97 0.629
$\omega\Xi(2101)$1.50 800 0.54 0.014 1.30 0.114
$\phi\Xi(2338)$1.57 800 0.75 0.012 1.80 0.077
@@
$a_{i}$ $\Lambda$ $|g_{i}|$ $\chi_{i}$ $|g_{i}|$ $\chi_{i}$
$\rho\Xi(2089)$2.19 1440 3.48 0.323 0.16 0.001
$\bar{K}^{*}\Lambda(2010)$2.00 1300 2.07 0.131 0.66 0.035
$\bar{K}^{*}\Sigma(2087)$1.87 1140 1.16 0.033 3.15 0.509
$\omega\Xi(2101)$1.50 800 0.41 0.004 1.43 0.100
$\phi\Xi(2338)$1.80 990 0.56 0.005 1.98 0.100
@@
$\rho\Xi(2089)$1.74 990 2.88 0.335 1.43 0.096
$\bar{K}^{*}\Lambda(2010)$1.42 800 1.41 0.116 1.53 0.186
$\bar{K}^{*}\Sigma(2087)$1.82 1090 2.58 0.250 2.66 0.309
$\omega\Xi(2101)$1.50 800 0.41 0.006 1.45 0.092
$\phi\Xi(2338)$1.85 1040 0.56 0.006 2.00 0.081
@@
As it can be seen in Table 5, our lighter state couples strongly to the $\rho\Xi$ channel. Note that this state, although its mass is lower than all the thresholds, has a non-zero width. This is because we take into account the width of the $\rho$ and the $\bar{K}^{*}$ in the loop function. For the higher energy peak we can see that it strongly couples to the $\bar{K}^{*}\Sigma$ channel. In the PDG data one can find two states in this region, the $\Xi(1820)$ and the $\Xi(1950)$ (Tab. 4). The first state has a known spin-parity $J^{\pi}=\frac{3}{2}^{-}$, which is compatible with our model.
@@
Figure 7: Solid lines represent the sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB 2. The vertical coloured dashed lines represent the location of the thresholds of the different channels and the black one represents the mass position of the $\Xi(1820)$ and the $\Xi(1950)$.
@@
Similarly to what we did in PB case, we will modify our subtraction constants within a reasonable range trying to adjust our results to the experimental data. Modifying the values of the subtraction constants we build model VB 2, which produces two resonances at energies compatible with the experimental masses of $\Xi(1829)$ and $\Xi(1950)$ (see Fig. 7 and Table 5). We note that the values of the equivalent cut-offs for the subtractions constants employed are not larger than $\Lambda=1500~{}{\rm MeV}$, which is a reasonable upper limit value (Tab. 5). However, the widths of these obtained resonances are much smaller than experimental ones. Potentially this problem might be solved if we developed a model which would couple the VB channels with the PB ones, since this would open more light channels in which these resonances could decay, increasing their width. Such a coupling was studied in Garzon:2012np , and indeed it generated some increase of the width for the lower lying resonance.
@@
Figure 8: Solid lines represent the sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB 3. The vertical coloured dashed lines represent the location of the thresholds of the different channels and the black one represents the mass position of the $\Xi(1950)$.
@@
Another possibility to adjust our model to the experimental data is to generate both resonances in the energy region where the $\Xi(1950)$ appears. As described in the PDG PDG this state could be representing more than one $\Xi$ resonance, eventually merging into an apparent wide peak. With such a motivation we generated model VB 3, which, as can be seen in Fig. 8 and Table 5, produces two states close to the experimental position of the $\Xi(1950)$. Note that for this model we only need to modify three subtraction constants and the equivalent cut-off values are not larger than $1100~{}{\rm MeV}$ for all the channels. If our resonances were a bit wider, what can be achieved, for example, by coupling our VB channels with the PB ones, as discussed above, they would produce just one, but rather wide, peak in the experimental data, as it happens for $\Lambda(1405)$.
@@ III.2 Heavy sector
Figure 9: Solid lines represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained applying the dimensional regularisation scheme for the loop function. The vertical dashed lines represent the location of the thresholds of the different channels.
@@
Let us now discuss the results in the heavy sector, where there are only four channels. In a rather brief form these results have already been published in ref Marse-Valera:2022khy , while in this section we will show more results and more details of the calculations.
@@
We will start in a similar way as we did in the previous section. First of all, we set the regularisation scale to $\mu=1000~{}{\rm MeV}$ and the subtraction constants are evaluated for all the channels mapping the loop function to the value obtained for a cut-off of $\Lambda=800~{}{\rm MeV}$ at the threshold of each channel. In such a way four peaks are generated, as we can see in Fig. 9.
@@
Figure 10: The purple line represents the inverse of the diagonal term of the potential, the green line shows the loop function of the same channel, computed using a dimensional regularisation scheme, and the blue line is similar to the green one, but computed using a cut-off scheme. In (a) we have the $\bar{D}_{s}\Xi_{c}$ channel, in (b) the $\bar{D}_{s}\Xi_{c}^{\prime}$ and in (c) the $\bar{D}\Omega_{c}$.
@@
Analyzing these results wee observe that the first three peaks couple very strongly to $\bar{D}_{s}~{}\Xi^{\prime}_{c}$ and/or to $\bar{D}_{s}~{}\Xi_{c}$, what is very suspicious, since the diagonal elements of the potential for these channels are repulsive ($C_{ii}<0$, see Tab. 1). This might be an indication that these peaks/poles are fake; these might be mathematical artifacts of the dimensional regularization scheme, that have been already seen before. To illustrate the problem, we present in Figs. 10 the inverse of the diagonal potential and the real part of the loop function calculated both using the dimensional regularisation and the cut-off methods for three heaviest channels. Note that according to Eq. (12), a pole is generated when $VG\sim 1$. As we can see in Figs. 10 (a) and (b) the inverse of the potential and the real part of the dimensional regularized loop function almost intersect in a region where the potential is repulsive. Thus, these states, which mostly couple to $\bar{D}\Xi_{c}$ and $\bar{D}\Xi_{c}^{\prime}$, are only mathematical solutions of our equation, but do not represent any physical state. Fig. 10 also shows that to avoid such fake poles we should use the cut-off scheme to compute the loop functions.
@@
Figure 11: Solid lines represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, obtained using the cut-off model to compute the loop function. The vertical dashed lines represent the location of the thresholds of the different channels.
@@
As can be seen in Fig. 11, using a cut-off model with $\Lambda=800~{}{\rm MeV}$, we make the fake poles disappear and, as expected, only the pole at higher energy remains. This one strongly couples to the $\bar{D}\Omega_{c}$ and $\bar{D}_{s}\Xi_{c}^{\prime}$ channels (see Table 6). This state, which would be a pentaquark with double strangeness ($P_{\Psi_{ss}}^{\Xi}$) has a well defined spin-parity $J^{\pi}=\frac{1}{2}^{-}$ as it is generated from the PB interaction in s-wave.
@@
It is very interesting to look for the origin of this resonance. Checking Table 1 we can see that there is only one attractive diagonal term, for $\bar{D}\Omega_{c}$ channel, and although it is positive (i.e. attractive in our parameterization) itis still rather small small $(C_{99}=1/9)$. Such an attraction is not sufficient to generate a resonance state (what was check performing uncoupled calculations). Thus, this resonance is generated in a rather unusual way, via the non-diagonal terms of the interaction. The $\bar{D}\Omega_{c}-\bar{D}_{s}\Xi_{c}^{\prime}$ interaction, $C_{98}=-\sqrt{2}$, is strong enough to produce a bound state via the coupled-channels. To check this point, we added a $\gamma$ factor, $C_{89}\to C_{89}^{\prime}=-\gamma\sqrt{2}$, which can be varied to see the effect of this off-diagonal term in details. In Fig. 12 we can clearly see the effect of this term - for smaller values of $C_{89}^{\prime}$ ($\gamma\sim 0.5$) the resonance state is not generated.
Figure 12: Solid lines represent the sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for $i=\bar{D}\Omega_{c}$ and for different $\gamma$’s ($C_{89}\to C_{89}^{\prime}=\gamma C_{89}$). Loop function is calculated with a cut-off scheme as in Fig. 11. The vertical dashed lines represent the location of the thresholds of the different channels.
Figure 13: Solid lines represent the sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained using the cut-off scheme to compute the loop function. The vertical dashed lines represent the location of the thresholds of the different channels.
We now present our results in the heavy VB sector. In the similar way as in the PB sector, the use of dimension regularization for the loop function leads to some fake nonphysical poles, which have been eliminated by the use of the cut-off method. Finally, only one resonance state remains, which couples strongly to $\bar{D}^{*}\Omega_{c}$ and $\bar{D}_{s}^{*}\Xi_{c}^{\prime}$, see Fig. 13 and Table 7. Similarly to the PB sector this state can be formed due to the large value of $C_{89}$, what is illustrates in Fig. 14. Note that this state is degenerate in spin-parity, can be either $J^{P}=\frac{1}{2}^{-}$ or $J^{P}=\frac{3}{2}^{-}$, since we have a vector-meson $(J^{P}=1^{-})$ interacting with a baryon $(J^{P}=\frac{1}{2}^{+})$ in s-wave.
@@
Figure 14: Solid lines represent the sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$, for $i=\bar{D}^{*}\Omega_{c}$ and for different $\gamma$’s ($C_{89}\to C_{89}^{\prime}=\gamma C_{89}$). Loop function is calculated with a cut-off scheme as in Fig. 13. The vertical dashed lines represent the location of the thresholds of the different channels.
We also have checked that the variations of cut off value, of $\kappa_{c}$ and $\kappa_{cc}$ parameters (related to explicit violation of SU(4) symmetry) lead to some variations of the resonance position and properties, but the appearance of a pole is a robust outcome in all calculations both for PB and VB cases Marse-Valera:2022khy .
Thus we have shown that in a simple model with realistic parameters we have been able to dynamically generate three pentaquarks with doubly strangeness and hidden charm. Shortly after our first results have been published Marse-Valera:2022khy , another study performed with a similar model appeared Roca:2024nsi , where similar pentaquarks have been generated in the same mass region, but with much smaller widths. The main reason for this differences, as it is explained by the authors of Roca:2024nsi , is the $-1/\sqrt{3}$ factor that appears in their $C{ij}$ constants for the $\eta_{c}\Xi$ and $J/\psi\Xi$ channels. We checked that by adopting our matrix to the values Roca:2024nsi , our model generates resonance practically at the same place, but with a smaller width of about $22~{}{\rm MeV}$. Our states are still generated at a bit higher energies, what makes it easier for them to decay into the $\eta_{c}\Xi$ or the $J/\psi\Xi$ channels increasing their width.

### v2

@@ III.1 Light sector
We first discuss the results of the light channels of the pseudoscalar meson baryon interaction, where we employ the dimensional regularisation scheme to compute the loop function. The values of the unknown subtraction constants are determined by imposing the loop function to coincide, for a regularisation scale of $\mu=1000~{}{\rm MeV}$, with the cut-off loop function with $\Lambda=800~{}{\rm MeV}$ at the corresponding threshold. [^1] Note that, in this light sector, the cut-off loop function employed to determine the subtraction constants is not that of Eq. (15), but the alternative expression which is obtained when only the positive energy term of the baryon propagator is retained. This is the expression employed e.g. in [3]. We chose this value as it corresponds to the mass of the interchanged vector meson on the t-channel diagram which integrated out when we take the $t\to 0$ limit. We will refer to this procedure as Model PB1.
@@
Figure 3: Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{3j}|$, obtained in Model PB1, as a function of the center-of-mass energy. The solid line shows the results when the heavy channels are neglected and the dashed one presents the results using all nine channels. The vertical coloured dashed lines represent the location of the thresholds of the different channels.
@@
Figure 4: (a): Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for a fixed channel $i$, obtained in Model PB1 as a function of the center-of-mass energy. (b): Same as (a) but showing the region of the second pole with more details. The vertical dashed lines represent the location of the different channel thresholds.
@@
$a_{i}$ $\Lambda~{}~{}~{}~{}$ ${}~{}~{}~{}|g_{i}|~{}~{}~{}$ ${}~{}~{}\chi_{i}~{}~{}~{}$ ${}~{}~{}|g_{i}|~{}~{}~{}$    $\chi_{i}$
$\pi\Xi(1456)$ $-1.51$ $800$ $2.44$ $0.428$ $0.12$ $0.001$
$\bar{K}\Lambda(1611)$ $-1.32$ $800$ $1.91$ $0.169$ $0.27$ $0.006$
$\bar{K}\Sigma(1689)$ $-1.27$ $800$ $0.71$ $0.016$ $1.54$ $0.815$
$\eta\Xi(1866)$ $-1.47$ $800$ $0.49$ $0.004$ $1.06$ $0.029$
@@
We can observe that the lowest generated resonance couples strongly to the $\pi\Xi$ and $\bar{K}\Lambda$ channels, which are open for decay, what actually explains its rather large width. In the PDG [7] one can find a state with similar mass, $\Xi(1620)$, which was seen decaying into the $\pi\Xi$ channel (see Table 4). Even if our state is $50~{}{\rm MeV}$ lighter and its width is larger, we can see the potential of our model to generate a state in this region.
As it can also be seen in Table 3 our heavier state mostly couples to the $\bar{K}\Sigma$ and $\eta\Xi$ channels and it has a mass of $1685~{}{\rm MeV}$. According to the PDG there exists a $\Xi(1690)$ state at $1690\pm 10~{}{\rm MeV}$ with a width of $20\pm 15~{}{\rm MeV}$ (see Table 4), which is compatible with our result. If this state was moved to a bit higher energy, the $\bar{K}\Sigma$ channel would be open for it to decay, and this would naturally lead to an increase of the width.
@@
Looking at the results, it is natural to think that we can modify the values of the subtraction constants within a reasonable range in order to accommodate the generated states to the experimental data. Therefore, we relax the condition of the loop function matching the cut-off one with $\Lambda=800~{}{\rm MeV}$ at the corresponding threshold [33]. By doing that we define a new model, called PB2. The results of new calculations within Model PB2 are shown in Fig. 5 and in Table 3. We can see that the lighter resonance now appears at higher mass and it is much wider, while the heavier resonance is generated with a similar mass and with a slightly larger width. Looking at Table 4 we can see that now the mass of the light state is closer to the experimental energy of the $\Xi(1620)$. In this model, none of the equivalent cut-off values for the subtractions constants employed are smaller than $\Lambda=630~{}{\rm MeV}$, which is a reasonable lower limit value.
Figure 5: Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for a fixed channel $i$, obtained in Model PB2 as a function of the center-of-mass energy. The vertical dashed lines represent the location of the different channel thresholds.
Reviewing the literature [41, 2, 3, 42, 43, 44, 29, 45, 46] one can note that all the models based on just a leading interaction term in the potential, the so called Weinberg-Tomozawa term, see Eq. (6), can dynamically generate two poles, the features of which cannot reproduce the experimental masses and widths of the $\Xi(1620)$ and the $\Xi(1690)$ at the same time. Furthermore, the width of the $\Xi(1690)$ remains always rather small, around a few MeV. Consistently, our calculations also show all these features. In some cases the $\Xi(1620)$ resonance can be reproduced very well [29, 45, 46], but only allowing some of the subtracting constants to take unnatural-sized values.
These limitations can be overcome if one also takes into consideration the s- and u-channel Born diagrams and the NLO contribution, as it was done in [30, 47, 32], where the properties of both $\Xi(1620)$ and $\Xi(1690)$ resonances are well reproduced by the model.
@@
Figure 6: Sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB1 as a function of the center-of-mass energy. The vertical coloured dashed lines represent the location of the different channel thresholds and the black dot-dashed one denotes the mass position of the $\Xi(1950)$.
@@
$\rho\Xi(2089)$ $-1.50$ 800 3.39 0.610 0.24 0.004
$\bar{K}^{*}\Lambda(2010)$ $-1.42$ 800 1.82 0.358 0.52 0.0490
$\bar{K}^{*}\Sigma(2087)$ $-1.46$ 800 0.98 0.047 2.97 0.629
$\omega\Xi(2101)$ $-1.50$ 800 0.54 0.014 1.30 0.114
$\phi\Xi(2338)$ $-1.57$ 800 0.75 0.012 1.80 0.077
@@
${}~{}a_{i}~{}~{}$ ${}~{}\Lambda~{}~{}$ ${}~{}~{}|g_{i}|~{}~{}~{}$ ${}~{}~{}\chi_{i}~{}~{}~{}$ ${}~{}~{}|g_{i}|~{}~{}~{}$ ${}~{}~{}\chi_{i}~{}~{}~{}$
$\rho\Xi(2089)$ $-2.19$ 1440 3.48 0.323 0.16 0.001
$\bar{K}^{*}\Lambda(2010)$ $-2.00$ 1300 2.07 0.131 0.66 0.035
$\bar{K}^{*}\Sigma(2087)$ $-1.87$ 1140 1.16 0.033 3.15 0.509
$\omega\Xi(2101)$ $-1.50$ 800 0.41 0.004 1.43 0.100
$\phi\Xi(2338)$ $-1.80$ 990 0.56 0.005 1.98 0.100
@@
$\rho\Xi(2089)$ $-1.74$ 990 2.88 0.335 1.43 0.096
$\bar{K}^{*}\Lambda(2010)$ $-1.42$ 800 1.41 0.116 1.53 0.186
$\bar{K}^{*}\Sigma(2087)$ $-1.82$ 1090 2.58 0.250 2.66 0.309
$\omega\Xi(2101)$ $-1.50$ 800 0.41 0.006 1.45 0.092
$\phi\Xi(2338)$ $-1.85$ 1040 0.56 0.006 2.00 0.081
@@
As it can be seen in Table 5, the lighter state couples strongly to the $\rho\Xi$ channel. Note that, although the mass of this state is lower than all the channel thresholds, it has a non-zero width. This is because we take into account the width of the $\rho$ and the $\bar{K}^{*}$ in the loop function. We can see that the higher energy state couples strongly to the $\bar{K}^{*}\Sigma$ channel. In the PDG compilation one finds two states in this region, $\Xi(1820)$ and $\Xi(1950)$ (see Table 4). The first state has a known spin-parity $J^{\pi}=\frac{3}{2}^{-}$, which is compatible with our model.
@@
Figure 7: Sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB2 as a function of the center-of-mass energy. The vertical black dot-dashed lines represent the mass position of the $\Xi(1820)$ and the $\Xi(1950)$.
@@
Similarly to what we did in PB case, we will modify our subtraction constants within a reasonable range trying to adjust our results to the experimental data. In this way, we build model VB2, which produces two resonances at energies compatible with the experimental masses of $\Xi(1829)$ and $\Xi(1950)$ (see Fig. 7 and Table 5). We note that the values of the equivalent cut-offs for the subtractions constants employed are not larger than $\Lambda=1500~{}{\rm MeV}$, which is a reasonable upper limit value. However, the widths of these obtained resonances are much smaller than the experimental ones. This problem might potentially be solved if we implemented the coupling of the VB channels with the PB ones, as this would open more light channels in which these resonances could decay, thereby increasing their width. Such a coupling was studied in [48], via pseudoscalar exchange plus a contact (or Kroll-Ruderman) term (the latter turning out to be more important), and indeed it generated a moderate increase of the width for the lower lying resonance.
@@
Figure 8: Sum over all j channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained in Model VB3. The vertical coloured dashed lines represent the location of the thresholds of the different channels and the black one represents the mass position of the $\Xi(1950)$.
@@
Another possibility to adjust our model to the experimental data is to generate both resonances in the energy region where the $\Xi(1950)$ appears. As described in the PDG [7] this state could be representing more than one $\Xi$ resonance, eventually merging into an apparent wide peak. With such a motivation we generated model VB3, which, as can be seen in Fig. 8 and Table 5, produces two states close to the experimental position of the $\Xi(1950)$. Note that this model required the modification of only three subtraction constants and the equivalent cut-off values are not larger than $1100~{}{\rm MeV}$ for all the channels. If our resonances were a bit wider, which could be achieved by coupling our VB channels with the PB ones, as discussed above, they would produce just one, but rather wide, peak in the spectrum.
@@ III.2 Heavy sector
Figure 9: Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained with the dimensional regularization scheme for the loop function, as a function of the center-of-mass energy. The vertical dashed lines represent the location of the different channel thresholds.
@@
Let us now discuss the results in the heavy sector, where there are only four channels. Some results have already been published in Ref. [20] in a rather brief form. In this section we will provide more details of the calculations and and some additional results.
@@
Similarly to the previous section, we first set the regularization scale to $\mu=1000~{}{\rm MeV}$ and the subtraction constants are evaluated by mapping, at the threshold of each channel, the dimensional regularized loop function to the value obtained with the cut-off expression employing $\Lambda=800~{}{\rm MeV}$. In such a way four peaks are generated, as we can see in Fig. 9.
@@
Figure 10: The purple line represents the inverse of the diagonal term of the potential, the green line shows the loop function of the same channel, computed using a dimensional regularization scheme, and the blue line is similar to the green one, but computed using a cut-off scheme. In (a), (b) and (c) we show the $\bar{D}_{s}\Xi_{c}$, $\bar{D}_{s}\Xi_{c}^{\prime}$ and $\bar{D}\Omega_{c}$ channel, respectively.
@@
Analyzing these results wee observe that the first three peaks couple very strongly to $\bar{D}_{s}~{}\Xi^{\prime}_{c}$ and/or to $\bar{D}_{s}~{}\Xi_{c}$, which is surprising since the diagonal elements of the potential for these channels are repulsive ($C_{ii}<0$ in Table 1). This might be an indication that these peaks might be mathematical artifacts of the dimensional regularization scheme. To illustrate the problem, we present in Figs. 10 the inverse of the diagonal potential and the real part of the loop function, the latter calculated using both the dimensional regularization and the cut-off schemes, for the three heavier channels. Note that, according to Eq. (12), a pole is generated when $VG\sim 1$. As we can see in Figs. 10 (a) and (b) the inverse of the potential and the real part of the dimensional regularized loop function almost intersect in a region where the potential is repulsive. Thus, these states, which mostly couple to $\bar{D}\Xi_{c}$ and $\bar{D}\Xi_{c}^{\prime}$, are only mathematical solutions of our equation, but do not represent any physical state. Fig. 10 also shows such fake poles could be avoided by employing the cut-off scheme to compute the loop functions.
@@
Figure 11: Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, obtained with the cut-off model to compute the loop function, as a function of the center-of-mass energy. The vertical dashed lines represent the location of the different channel thresholds.
@@
As can be seen in Fig. 11, the fake poles disappear when we use a cut-off scheme with $\Lambda=800~{}{\rm MeV}$ and, as expected, only the pole at higher energy remains. This one strongly couples to the $\bar{D}\Omega_{c}$ and $\bar{D}_{s}\Xi_{c}^{\prime}$ channels (see Table 6). This state, which would be a pentaquark with double strangeness ($P_{\Psi_{ss}}^{\Xi}$) has a well defined spin-parity $J^{\pi}=\frac{1}{2}^{-}$ as it is generated from the PB interaction in s-wave.
@@
It is very interesting to look for the origin of this resonance. Checking Table 1 we can see that only diagonal $\bar{D}\Omega_{c}$ term is positive (i.e. attractive in our notation) although with a rather small strength $(C_{99}=1/9)$. Such an attraction is not sufficient by itself to give rise to a bound state, a fact that we checked upon performing an uncoupled calculation. Thus, as already noted in our previous work [20], this resonance is generated in a rather unusual way from the non-diagonal terms of the interaction. Indeed, the $\bar{D}\Omega_{c}-\bar{D}_{s}\Xi_{c}^{\prime}$ interaction is strong enough ($C_{98}=-\sqrt{2}$), to produce a bound state via the coupled channels. To check this point, we added a $\gamma$ factor to this coupling, namely $C_{89}\to C_{89}^{\prime}=-\gamma\sqrt{2}$, which we varied to study the effect of this off-diagonal term in detail, as can be seen in Fig. 12. We observe that the resonance is not generated for values of $\gamma\lesssim 0.5$.
Figure 12: Sum over all $j$ channels of the modulus of the PB scattering amplitude, $|T_{ij}|$, for $i=\bar{D}\Omega_{c}$ and for values of $\gamma$ ($C_{89}\to C_{89}^{\prime}=\gamma C_{89}$), as a function of the center-of-mass energy. The loop function is calculated with a cut-off scheme as in Fig. 11. The vertical dashed lines represent the location of the different channel thresholds.
Figure 13: Sum over all $j$ channels of the modulus of the VB scattering amplitude, $|T_{ij}|$ for a fixed channel $i$, obtained using the cut-off scheme to compute the loop function, as a function of the center-of-mass energy. The vertical dashed lines represent the location of the different channel thresholds.
We now present our results in the heavy VB sector. In the similar way as in the PB sector, the use of dimension regularization for the loop function leads to some fake nonphysical poles, which have been eliminated by the use of the cut-off method. Finally, only one resonance state remains, which couples strongly to $\bar{D}^{*}\Omega_{c}$ and $\bar{D}_{s}^{*}\Xi_{c}^{\prime}$, see Fig. 13 and Table 7. Similarly to the PB sector this state can be formed due to the large value of the $C_{89}$ coefficient. Note that this state is degenerate in spin-parity, can be either $J^{P}=\frac{1}{2}^{-}$ or $J^{P}=\frac{3}{2}^{-}$, since we have a vector-meson $(J^{P}=1^{-})$ interacting with a baryon $(J^{P}=\frac{1}{2}^{+})$ in s-wave.
@@
We also have checked that the variations of cut off value, of $\kappa_{c}$ and $\kappa_{cc}$ parameters (related to explicit violation of SU(4) symmetry) lead to some variations of the resonance position and properties, but the appearance of a pole is a robust outcome in all calculations both for PB and VB cases [20].
Thus we have shown that a simple model with realistic parameters does generate three pentaquarks with double strangeness and hidden charm. Shortly after our first results were published [20], another study performed with a similar model [22] predicted pentaquarks in the same mass region, but with much smaller widths. As explained in Ref. [22], the main reason for this difference is the relative factor of $-1/\sqrt{3}$ factor multiplying their $\bar{D}_{s}\Xi_{c}^{\prime}\to\eta_{c}\Xi$ and $\bar{D}\Omega_{c}\to\eta_{c}\Xi$ transitions. Indeed, we have checked that implementing theses factors our model generates resonances practically at the same place as before but with a smaller width of about $22~{}{\rm MeV}$. This width is still about twice as that found in [22] because our model also contemplates the additional $\bar{D}_{s}\Xi_{c}$ channel with a sizable coupling to $\eta_{c}\Xi$, the primary final decaying state.